Consider the following information for Evenflow Power Co.,

Debt: 3,500 8 percent coupon bonds outstanding, $1,000 par value, 22 years to maturity,selling for 102 percent of par; the bonds make semiannual payments.
Common stock: 84,000 shares outstanding, selling for $57 per share; the beta is 1.18.
Preferred stock: 11,500 shares of 7 percent preferred stock outstanding (note: multiply this percentage in decimal format times 100 to get the dividend), currently selling for $105 per share.
Market: 8.5 percent market risk premium and 6.5 percent risk-free rate.

Assume the company's tax rate is 32 percent.

Required:

Find the WACC.(Do not round your intermediate calculations.)
Answer #1
Calculation of cost of debt

Price =102%*$1000 =1020
Let cost of debt be r
Semiannual coupon payment = 8%*1000/2= 40
1020 = 40/(1+r/2) + 40/(1+r/2)^2 + 40/(1+r/2)^2 ...............10 40/(1+r/2)^44

r =7.80%
Total Debt =1020*3,500 = $3,570,000

Calculation of cost of equity
Re = 6.5% +1.18**8.5% =16.53%
Total Equity = 84,000*$57=$4,788,000

Calculation of cost of preferred stock
Dividend = 7%*100 =7
Cost of Preferred stock = 7/105 =6.67%
Value of Preferred Stock =$105*11,500 =$ 1,207,500

WACC = (D/V)*Rd*(1-T) + (E/V)*Re + (P/V)*Rp

WACC = (16.53%*$4,788,000 +6.67%*$1,207,500 + 7.80%*$3,570,000*(1-.32))/($4,788,000 + $1,207,500+ $3,570,000)
= 11.10%